Page 2 11) The fastest measured pitched baseball left the pitchers hand at a speed of 45 m/s. If the pitcher was in contact with the ball over a distance of 1.5 m and produced constant acceleration a) what acceleration did he give the ball, and b) how much time did it take him to pitch it? 12) A car accelerates from 20 m/s to 50 m/s in 3 s. What is the cars acceleration? 13) A car accelerates from 20 m/s to 50 m/s over a distance of 100 m. What is the cars acceleration? 14) A small block has constant acceleration as it slides down a frictionless incline. The block is released from rest at the top of the incline, and its speed after it has traveled 6.8 m to the bottom of the incline is 3.8 m/s. What is the speed of the block when it is 3.4 m from the top of the incline? 4 Previous
Answer #1

11) a = v2 - u2 / 2 s

= 452 / (2 * 1.5)

acceleration = 675 m/s2

b) t = v - u / a = 45 / 675

time required = 0.066 s

12) a = v - u / t

= (50 -20) / 3

cars acceleration = 10 m/s2